OpenStudy (anonymous):

In set-builder notation, how do you write the solutions of the inequality? -4 - z *>=* -2 A) {x|x *>=* -2} B) {x|x *>=* 2} C) {x|x *<=* -2} D) {x|x *<=* 2}

OpenStudy (anonymous):

\[-4-z \ge-2,-z \ge -2+4,-z \ge 2,z \le -2,z=\left\{ x:x \le-2 \right\}\]

OpenStudy (anonymous):

huh ?

OpenStudy (dtan5457):

@surjithayer what he/she did here was isolate z. that's the hard part. {x:x≤−2} means the set of all x such that x is <=-2

OpenStudy (anonymous):

C ?

OpenStudy (dtan5457):

yep

OpenStudy (anonymous):

THANKS !!!